An air plane travels 280 m through

the runway before taking off. It

starts from rest, moves. with

constant acceleration and takes off

after 8 s. What is its speed when it

takes off?


1
Expert's answer
2021-02-25T11:23:23-0500

The distance traveled by air plane before taking off

"d=\frac{v_f+v_i}{2}t"

Hence, the final velocity

"v_f=2d/t-v_i=2*280/8-0=70\;\rm m/s"